OpenStudy (anonymous):

3 {sqrt}y + 6 {sqrt}y

OpenStudy (anonymous):

\[3\sqrt{y}+6\sqrt{y}\]

OpenStudy (anonymous):

simplify by collecting radical terms

OpenStudy (cwrw238):

what is 3x + 6x?

OpenStudy (anonymous):

9x

OpenStudy (cwrw238):

ok so let sqrty = x

OpenStudy (anonymous):

\[9\sqrt{y}\]

OpenStudy (cwrw238):

u can treat radicals like variables

OpenStudy (cwrw238):

yup

OpenStudy (anonymous):

ok thanks

OpenStudy (cwrw238):

yw